For every function f(x) that is smoothly continuous on (a,b], the average rate of change of f(x) from x = a to x = b, approaches the instantaneous rate of change at x = a as x approaches a from the left.


Please explain this with a graph


1
Expert's answer
2021-12-17T13:48:55-0500

Average rate of change:

"(f(b)-f(a))\/(b-a)=f'(a)"

since "x=a^+" , which is in the interval [a,b].


If b=a+h where h is very small,

"f'(a)=(f(a+h)-f(a))/h\ as\ h→0"

by definition of derivative (instantaneous rate of change).